OpenStudy (anonymous):

Will someone please finish helping me...been asking for weeks...really need the right answer... I am seriously having issues trying to figure out two of these simple answers....please help Of 50 employees of a store located in downtown Boston, 18 people take the subway to work, 12 take the bus, and 7 take both the subway and the bus. How many employees (a) Take the subway to the bus to work? (b) Take only the bus to work? (c) Take either the bus or the subway to work? (d) Get to work by some other means? I already figured that (b) = 12-7=5 I already figured that (c) = 11+5=16 I cannot get (a)

OpenStudy (anonymous):

I just DO NOT GET IT YET...

OpenStudy (anonymous):

I think I got A....11+5 = 15..............

OpenStudy (anonymous):

16 i mean

OpenStudy (anonymous):

a.18:12 it is asking for a ratio of people that take the subway to people that take the bus to work. b.12 people take the bus . that information is given in the passage. c.18+12=30 people because 18 take the subway and 12 take the bus you just add d. 2o because you already know 30 take a bus or ride the subway and you just subtract 50-30=20 people left